OpenStudy (anonymous):

Integral from 1 to 2 of ( v (e-e inverse) dv

OpenStudy (anonymous):

\[\int\limits_{1}^{2}v (e-e^-1)\]

OpenStudy (anonymous):

easy

OpenStudy (amistre64):

e is a constant; and so is 1/e

OpenStudy (anonymous):

2 ( e-e^-1 )

OpenStudy (anonymous):

wait thats wrong lol

OpenStudy (anonymous):

take the constant out

OpenStudy (amistre64):

might as well ask what is the integral of 4v dv :)

OpenStudy (anonymous):

integral v from 2 to 1

OpenStudy (anonymous):

(1/2) [ 4-1] = 3/2 so \[\frac{3}{2} (e - e ^{-1})\]

OpenStudy (anonymous):

^ answer

OpenStudy (amistre64):

\[F(v)=\frac{v^2(e-e^{-1})}{2} |_{1}^{2}\]

OpenStudy (amistre64):

e - (1/e) = (e^2 - 1)/e right?

OpenStudy (anonymous):

So don't treat v as a constant

OpenStudy (anonymous):

its dv on the end

OpenStudy (amistre64):

v is the variable here

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So which answer is correct

OpenStudy (amistre64):

4.7008047745751485263396882938069 -1.1752011936437871315849220734517 -------------------------------------- 3.525603580931361394754766220355 is what i get :)

OpenStudy (anonymous):

:|

OpenStudy (amistre64):

F(2) - F(1)

OpenStudy (anonymous):

stuff like that is what puts people off maths ^ :|

OpenStudy (anonymous):

no one cares about the numbers, especially to 20 decimal places , the final number doesnt matter , its how you go the number that matters Whenever you have e's or \[\pi\] in the answer then the answer should be given as exact value

OpenStudy (anonymous):

are you in kindergarten amistre lol

OpenStudy (amistre64):

i gave the exact and the approximate; or so i thought....

OpenStudy (amistre64):

i even gave a reason of "how" it is obtained.... right?

OpenStudy (anonymous):

your like the guys that memorise pi to 1000's of demical places

OpenStudy (amistre64):

lol.... 7 digits on pi :) e = 2.71828182845905... is that best i got with that one :)

OpenStudy (anonymous):

never give an answer to 20decimal places, makes you lokok like an idiot, 2 ( maybe 3 ) max , no one cares bout the rest. When a person that is new to maths sees something really long and unneccessary like that it is sure to put them off

OpenStudy (anonymous):

*look like .....

OpenStudy (amistre64):

i dont have to include decimals to look like an idiot; im a pro at that regardless :) and i figured that if they are asking calculus questions then they arent all that new to math...